Complete the following fusion reaction: 4He+3He⟶2H+−−−−−−.

Express your answer as a complete nuclear equation.

nothing

SubmitRequest Answer
Answer #1

Atomic number of missing particle,

n = (2 + 2) - 1 = 3

atomic mass of missing particle,

A.m = 4 + 3 - 2 = 5

so, particle is 3Li5 .

fusion reaction:

2He4 + 2He31H2 + 3Li5